A random sample of 8 size AA batteries for toys yielded a mean lifespan of 3.65 hours with standard deviation, 0.76 hours. In a normal probability plot of battery lifespans, all of the points fell between the curved lines.

Find the margin of error for a 99% CI.
Answer #1

Solution :

Given that,

t /2,df = 3.499

Margin of error = E = t/2,df * (s /n)

= 3.499 * (0.76 / 8)

Margin of error = E = 0.94